OpenStudy (anonymous):

find all of the fourth roots of 81(cos(3pi/8) + i sin (3pi/8) express your answer in polar (r(cos theta+ i sin theta)) form

OpenStudy (anonymous):

r = 81^1/4=3

OpenStudy (anonymous):

i will be with you in a moment just got a call...

OpenStudy (anonymous):

recall ... if \[w = r(\cos(\theta)+isin(\theta))\] is a complex number then in order to obtain the nth root that yields to \[\sqrt[n]{w}=r^{1/n}[\cos(\frac{\theta+2k \pi}{n})+ isin(\frac{\theta+2k \pi}{n})]\] where k=1,2,3...n-1

OpenStudy (anonymous):

that would be the answer or would I have to do something else?

OpenStudy (anonymous):

sorry k=0,1,2,...n-1

OpenStudy (anonymous):

that would give you the exact answer to your question with no hassle...

OpenStudy (anonymous):

but i will help you solve it don't worry

OpenStudy (anonymous):

choose n=4, then k=n-1, so k=3

OpenStudy (anonymous):

your theta=3pi/8

OpenStudy (anonymous):

and you are done, so give the values k=0,1,2, and 3

OpenStudy (anonymous):

if you can't follow me then tell me to solve the whole thing ...

OpenStudy (anonymous):

do I need to plug in something into the first equation you gave ?

OpenStudy (anonymous):

yes i told you plug k=0, then k=1, then k=2, then k=3, n =4 always

OpenStudy (anonymous):

n=4 because you want to find the 4th root

OpenStudy (anonymous):

wheredo I plug in the k variables?

OpenStudy (anonymous):

\[\sqrt[n]{w}=r^{1/n}[\cos(\frac{\theta+2k \pi}{n})+ isin(\frac{\theta+2k \pi}{n})]\]

OpenStudy (anonymous):

let me solve the first root for you as an example then you may generate the rest...ok!

OpenStudy (anonymous):

\[\sqrt[4]{w}=r^{1/4}[\cos(\frac{\theta+2(0) \pi}{4})+ isin(\frac{\theta+2(0) \pi}{4})]\]

OpenStudy (anonymous):

now notice for every n i substituted 4, and for every k i pluged in 0

OpenStudy (anonymous):

and thata would be 3pi/8?

OpenStudy (anonymous):

let me simplify this

OpenStudy (anonymous):

\[\sqrt[4]{w}=r^{1/4}[\cos(\frac{\frac{3\pi}{8}}{4}+\frac{2\pi (0)}{4})+ isin(\frac{\frac{3\pi}{8}}{4}+\frac{2\pi (0)}{4})]\]

OpenStudy (anonymous):

do you want me to simpliy this even more...

OpenStudy (anonymous):

yes please because I dont get what I would have to do next after this step

OpenStudy (anonymous):

\[\sqrt[4]{w}=r^{1/4}[\cos(\frac{3\pi}{32})+ isin(\frac{3\pi}{32})]\]

OpenStudy (anonymous):

how is that now?

OpenStudy (anonymous):

you see when k was zero we got rid of 2pik/4

OpenStudy (anonymous):

yes because it was zero.. so that would be the first root?

OpenStudy (anonymous):

the next root let k=1

OpenStudy (anonymous):

in this case 3pi/32+2pi/4

OpenStudy (anonymous):

all you worry about is what is inside the cos and sin

OpenStudy (anonymous):

how did you get 2pi/4?

OpenStudy (anonymous):

i got it from k=1, then 2(1)pi/4

OpenStudy (anonymous):

which is 2kpi/n

OpenStudy (anonymous):

3pi/32+2pi/4=19pi/32

OpenStudy (anonymous):

oh okay I got it now! Thank you so much!

OpenStudy (anonymous):

you are welcome but you need to find it till k=3

OpenStudy (anonymous):

yes and now I understand how to do it so I'll do it on my own Thank you

OpenStudy (anonymous):

this method is a secured method if you can memorize the formula, it will reduce the chance of making errors

OpenStudy (anonymous):

ok take care